OpenStudy (anonymous):

Find the slope and the y-intercept of a line. Write fractional answers in lowest terms. 7y+4x+3=5+4x

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Hmm.. you seem to have a 4x on both sides of your equation so you could cancel those....

OpenStudy (anonymous):

7y+4x+3=5+4x 7y=2 y=2/7 this is the line: y=2/7 parallel to x-axis the slope of that type of line =0

OpenStudy (anonymous):

Thanks!